# FinTSBridge: A New Evaluation Suite for Real-world Financial Prediction with Advanced Time Series Models

### v1

@@ Abstract
Despite the growing attention to time series forecasting in recent years, many studies have proposed various solutions to address the challenges encountered in time series prediction, aiming to improve forecasting performance. However, effectively applying these time series forecasting models to the field of financial asset pricing remains a challenging issue. There is still a need for a bridge to connect cutting-edge time series forecasting models with financial asset pricing. To bridge this gap, we have undertaken the following efforts: 1) We constructed three datasets from the financial domain; 2) We selected over ten time series forecasting models from recent studies and validated their performance in financial time series; 3) We developed new metrics, msIC and msIR, in addition to MSE and MAE, to showcase the time series correlation captured by the models; 4) We designed financial-specific tasks for these three datasets and assessed the practical performance and application potential of these forecasting models in important financial problems. We hope the developed new evaluation suite, FinTSBridge, can provide valuable insights into the effectiveness and robustness of advanced forecasting models in finanical domains. Code can be found in https://anonymous.4open.science/r/FinTSBridge-0D6C.
Time Series, Financial Dataset, Benchmark
@@ 1 Introduction
Time series forecasting has become increasingly crucial in the financial domain, playing a critical role in decision-making processes related to asset pricing, risk management, and algorithmic trading (Tang et al., 2022; Lim & Zohren, 2021; Masini et al., 2023). Accurate forecasts are essential for tasks such as stock index prediction, option pricing, and modeling cryptocurrency volatility, as they help optimize investment strategies and mitigate market risks (Li et al., 2022; Doshi et al., 2024). However, the non-stationary nature of financial markets-shaped by factors like geopolitical events and investor sentiment-adds significant complexity to these challenges (Zheng et al., 2023). Despite advancements in machine learning for time series analysis, transforming state-of-the-art models into actionable financial insights remains an ongoing obstacle (De Prado, 2018). For example, while models may achieve low mean squared error (MSE) on synthetic datasets, their performance often deteriorates in real-world conditions due to unaccounted market dynamics, such as price-volume interactions and multi-scale volatility.
Moreover, many studies focus on time series datasets with simplified statistical properties, such as electricity consumption or traffic flow, which typically exhibit strong stationarity and periodicity (Chen et al., 2024). Even commonly used financial datasets, such as historical exchange data, suffer from key limitations: daily resolution obscures intraday fluctuations, and limited variable coverage (e.g., the absence of derivatives metrics) fails to capture the full spectrum of multi-scale market interactions (Bao et al., 2024). Compounding these issues, traditional evaluation metrics like MSE and mean absolute error (MAE) prioritize point-wise accuracy while neglecting temporal correlations, a critical aspect in financial forecasting. For instance, our empirical results show that a naive model that predicts the last observed value can achieve competitive MSE but exhibits near-zero correlation with true market trends, making it ineffective for practical trading strategies. These shortcomings underscore the need for financial-specific benchmarks that reflect the complexity of real-world markets, along with evaluation frameworks that assess both predictive accuracy and economic utility.
@@ 2.1 Time-series Forecasting Models
Recent advances in time-series forecasting have focused on enhancing model capabilities through multi-scale decomposition, attention mechanism optimization, lightweight architectures, and representation learning innovations. Decomposition-based methods remain foundational for handling non-stationary signals. For example, Autoformer (Wu et al., 2021a) pioneers autocorrelation-driven periodicity detection, replacing traditional moving averages with adaptive decomposition, while FEDformer (Zhou et al., 2022) leverages Fourier-Wavelet hybrid spectral analysis to achieve multi-resolution frequency decomposition. Building on this, Non-stationary Transformers (Liu et al., 2022) introduce dynamic stationarization modules that learn to normalize non-stationary inputs before applying attention mechanisms, significantly improving robustness to distribution shifts. These approaches highlight the synergy between frequency-domain analysis and temporal modeling.
@@
To address computational bottlenecks in Transformers, researchers have developed sparse attention variants: Informer (Zhou et al., 2021) reduces quadratic complexity via probabilistic sparse attention with KL-divergence-based token selection, while Crossformer (Zhang & Yan, 2023) designs hierarchical cross-resolution attention to capture dependencies across temporal scales. Koopa (Liu et al., 2023a) bypasses attention entirely, proposing time-varying Koopman dynamic systems that model latent state evolution through linear operators. Concurrently, lightweight architectures challenge conventional wisdom-DLinear (Zeng et al., 2023) demonstrates that decoupled linear projections for trend and residual components can outperform complex models, while TSMixer (Chen et al., 2023) and TimeMixer (Wang et al., 2024a) adopt pure MLP-based architectures and hybrid frequency-time operators, respectively, to balance accuracy and computational cost.
@@
Representation learning breakthroughs further expand modeling capabilities. TimesNet (Wu et al., 2023) transforms 1D sequences into 2D temporal matrices via period-phase folding, enabling 2D convolutions to capture intra-period and inter-period patterns simultaneously. PatchTST (Nie et al., 2023) introduces channel-independent patching inspired by vision transformers, learning localized temporal embeddings through overlapping segments. iTransformer (Liu et al., 2024) inverts the traditional architecture by treating variates as tokens and time points as features, enhancing multivariate dependency modeling. MICN (Wang et al., 2023) employs multi-scale dilated convolutional pyramids to extract local periodic features hierarchically. Complementary techniques like RevIN (Kim et al., 2021) address distribution shifts via bidirectional instance normalization, and TiDE (Das et al., 2023) integrates temporal encoding with dense residual connections for efficient long-horizon forecasting. These innovations collectively advance three core principles: 1) Hybridization of frequency-temporal analysis, 2) Strategic simplification guided by signal processing theory, and 3) Systematic robustness enhancement through normalization and distribution alignment.
@@ 2.2 Financial task-related Studies
Recent advances in financial time series forecasting have increasingly adopted multi-horizon prediction frameworks to capture evolving market dynamics. Unlike traditional approaches that focus on single-step forecasting (e.g., predicting next-step prices or up-down trends as in (Ding et al., 2015; Abe & Nakayama, 2018; Kraus & Feuerriegel, 2017; Sun et al., 2023; Li et al., 2024)), recent time series methods like (Liu et al., 2024; Wu et al., 2023; Liu et al., 2023a) employ sequence-to-sequence architectures to predict multi-step trajectories of exchange time series, which helps to understand future temporal dynamics but also introduces challenges for prediction. Although large models and Reinforcement Learning methods are becoming increasingly popular in financial time series forecasting (Nie et al., 2024; Li et al., 2024; Zong et al., 2024), the potential of small models and supervised learning approaches in financial time series forecasting has not been fully explored. It is important to build a bridge between cutting-edge models and financial time series tasks.
@@ 2.3 The Predictability of Asset Prices
Stock price prediction is indeed challenging, but it is not entirely impossible. This is evidenced by the long-standing work of quantitative hedge fund institutions, which continuously mine predictive indicators and patterns from historical datasets to cope with the ever-changing market environment, achieving excess returns. Moreover, some studies based on the Efficient Market Hypothesis (EMH) suggest that the stock market is not necessarily in the semi-strong form or strong-form efficiency as previously thought, which would prevent future price movements from being forecasted based on available historical information (Miller et al., 1970; Malkiel, 2003). On the contrary, many markets often fall between semi-strong and weak-form efficiency (Lo & MacKinlay, 1988), and (Efficiency, 1993) has validated the momentum effect in the stock market, where stocks that performed well in the past may continue to perform well in the future, thereby challenging the weak-form efficient market hypothesis. (Fama & French, 1996) discusses the explanatory power of multi-factor models for asset pricing anomalies, indirectly indicating the level of market efficiency. Additionally, advancements in technology have significantly enhanced data processing capabilities, leading to improvements in the performance of predictive models. This development brings data processing power and predictive performance that were previously unavailable in models based solely on historical data (Brogaard & Zareei, 2023; Leippold et al., 2022; Gu et al., 2020; Feng et al., 2020; Barberis et al., 2005).
@@ 3 Dataset Curation
Current research on long-term time series forecasting predominantly focuses on eight mainstream datasets (Wu et al., 2023; Chen et al., 2023; Liu et al., 2024, 2023a; Zeng et al., 2023; Nie et al., 2023). Among these, five are related to electricity, while the remaining three pertain to weather, traffic, and exchange rates. Although exchange rate data falls under the financial domain, it is often overlooked compared to the other seven datasets or replaced by the ILI dataset from the illness area, due to the non-stationary and non-periodic characteristics of exchange rate data (Liu et al., 2023b), which makes it more challenging to predict than other time series data that exhibit significant periodicity and stationarity.
While some state-of-the-art time series models have demonstrated strong long-term forecasting capabilities on mainstream datasets, they lack the robustness needed to tackle the complexities of real-world time series problems (Tan et al., 2024; Bergmeir, 2024a, b, c). To better investigate these real-world challenges, we propose three financial time series datasets.
@@ 3.1 Data Sources
We have constructed three financial time series datasets: GSMI, OPTION, and BTCF, each representing different subfields of finance. The GSMI dataset includes 20 major indices from global stock markets, recording daily price and trading volume data for these indices over nearly 20 years, from 2005 to 2024. The OPTION dataset includes the CSI 300ETF options from the Chinese financial market, with variables related to risk for both call and put options. The BTCF dataset contains hourly frequency data for Bitcoin spot and perpetual contracts, helping to understand the spot-contract lag relationship and facilitate long-short trading strategies (Bulun, 2020; Narayanasamy et al., 2023). In Table 1, we present the statistical properties of these three financial time series datasets with more details.
Table 1: Summary of new datasets.
Dataset GSMI OPTION BTCF
Range (2000-01, 2024-12) (2024-04, 2024-12) (2020-01, 2024-11)
Variate 100 22 12
Samples 6533 37431 43014
Frequency Daily Minutely Hourly
Finance Aera Stock Indices Option Future & Spot
Predict Length {5,21,63,126} {5,21,63,126} {5,21,63,126}
@@ 3.2 Data Preprocessing Methods
For the trading volume series
$V_{0,t}=\{v_{0},v_{1},v_{2},...,v_{i},...,v_{t}\},$
the logarithmic transformation method is
@@ 3.3 Visualization of Preprocessing
Dataset Length Naive iTransformer DLinear FEDformer
MSE MAE MSE MAE MSE MAE MSE MAE
Exchange 96 0.081 0.196 0.086 0.206 0.088 0.218 0.148 0.278
192 0.167 0.289 0.177 0.299 0.176 0.315 0.271 0.315
336 0.306 0.398 0.331 0.417 0.313 0.427 0.460 0.427
720 0.810 0.676 0.847 0.691 0.839 0.695 1.195 0.695
ETTh2 96 0.281 0.369 0.297 0.349 0.333 0.387 0.358 0.397
192 0.339 0.404 0.380 0.400 0.477 0.476 0.429 0.439
336 0.372 0.424 0.428 0.432 0.594 0.541 0.496 0.487
720 0.439 0.463 0.427 0.445 0.831 0.657 0.463 0.474
ETTm2 96 0.202 0.310 0.180 0.264 0.193 0.292 0.203 0.287
192 0.235 0.337 0.250 0.309 0.284 0.362 0.269 0.328
336 0.270 0.361 0.311 0.348 0.369 0.427 0.325 0.366
720 0.335 0.401 0.412 0.407 0.554 0.522 0.421 0.415
Traffic 96 2.715 1.077 0.395 0.268 0.650 0.396 0.587 0.366
192 2.747 1.085 0.417 0.276 0.598 0.370 0.604 0.373
336 2.789 1.094 0.433 0.283 0.605 0.373 0.621 0.383
720 2.810 1.097 0.467 0.302 0.645 0.394 0.626 0.382
@@ 4 New Evaluation Metrics
While mainstream time series forecasting works in top conferences often adopt error-based methods, such as MSE and MAE, as evaluation metrics for model predictions, these metrics face significant challenges when applied to financial time series. A simple example is that a Naive model as in Table 2, which directly uses the last observed value of the input sequence as the forecast, achieves remarkably low prediction errors on the Exchange dataset (we provide detailed analysis in Section B.5). This raises critical questions about the evaluation of financial time series forecasting, suggesting that correlation-based metrics should be introduced alongside prediction errors, as they are crucial for real-world financial applications. Although traditional Information Coefficient and Information Ratio metrics (Treynor & Black, 1973; Grinold & Kahn, 2000) can measure temporal correlations in single-step univariate forecasting, they fail to assess multi-variable multi-step predictions. To address this limitation, we propose multi-step IC and multi-step IR, abbreviated as msIC and msIR, respectively.
@@ 5 Experiment
To bridge the gap between real-world financial time series data and cutting-edge time series models, we employ over 10 advanced time series models and conduct extensive experimental tests across three financial time series scenarios: Multivariate-to-Multivariate Forecasting, Multivariate-to-Univariate Forecasting, and Multivariate-to-Partial Forecasting. These time series models include: TimeMixer(Wang et al., 2024a), Koopa(Liu et al., 2023a), iTransformer(Liu et al., 2024), PSformer(Wang et al., 2024d), TiDE(Das et al., 2023), PatchTST(Nie et al., 2023), DLinear(Zeng et al., 2023), Stationary(Liu et al., 2022), TSMixer(Chen et al., 2023), TimesNet(Wu et al., 2023), FEDformer(Zhou et al., 2022), Autoformer(Wu et al., 2021a), Crossformer(Zhang & Yan, 2023), Transformer(Vaswani, 2017), Informer(Zhou et al., 2021), and a Naive model.
@@ 5.3 Multivariate-to-Partial Forecasting
Setup. Multivariate prediction of partial variables is not currently the mainstream experimental setup in time series forecasting. However, (Wang et al., 2024c) discusses the importance of this practical scenario. In this work, we set the closing prices of 20 indices in the GSMI dataset as the target variables for prediction. We evaluate the performance of the models on the GSMI dataset. Additionally, we construct a portfolio selection strategy and backtest the return performance while holding different numbers of indices simultaneously.

### v2

@@ Abstract
Despite the growing attention to time series forecasting in recent years, many studies have proposed various solutions to address the challenges encountered in time series prediction, aiming to improve forecasting performance. However, effectively applying these time series forecasting models to the field of financial asset pricing remains a challenging issue. There is still a need for a bridge to connect cutting-edge time series forecasting models with financial asset pricing. To bridge this gap, we have undertaken the following efforts: 1) We constructed three datasets from the financial domain; 2) We selected over ten time series forecasting models from recent studies and validated their performance in financial time series; 3) We developed new metrics, msIC and msIR, in addition to MSE and MAE, to showcase the time series correlation captured by the models; 4) We designed financial-specific tasks for these three datasets and assessed the practical performance and application potential of these forecasting models in important financial problems. We hope the developed new evaluation suite, FinTSBridge, can provide valuable insights into the effectiveness and robustness of advanced forecasting models in finanical domains.
@@ 1 Introduction
Time series forecasting has become increasingly crucial in the financial domain, playing a critical role in decision-making processes related to asset pricing, risk management, and algorithmic trading Tang et al. (2022); Lim & Zohren (2021); Masini et al. (2023). Accurate forecasts are essential for tasks such as stock index prediction, option pricing, and modeling cryptocurrency volatility, as they help optimize investment strategies and mitigate market risks Li et al. (2022); Doshi et al. (2024). However, the non-stationary nature of financial markets-shaped by factors like geopolitical events and investor sentiment-adds significant complexity to these challenges Zheng et al. (2023). Despite advancements in machine learning for time series analysis, transforming state-of-the-art models into actionable financial insights remains an ongoing obstacle De Prado (2018). For example, while models may achieve low mean squared error (MSE) on synthetic datasets, their performance often deteriorates in real-world conditions due to unaccounted market dynamics, such as price-volume interactions and multi-scale volatility.
Moreover, many studies focus on time series datasets with simplified statistical properties, such as electricity consumption or traffic flow, which typically exhibit strong stationarity and periodicity Chen et al. (2024). Even commonly used financial datasets, such as historical exchange data, suffer from key limitations: daily resolution obscures intraday fluctuations, and limited variable coverage (e.g., the absence of derivatives metrics) fails to capture the full spectrum of multi-scale market interactions Bao et al. (2024). Compounding these issues, traditional evaluation metrics like MSE and mean absolute error (MAE) prioritize point-wise accuracy while neglecting temporal correlations, a critical aspect in financial forecasting. For instance, our empirical results show that a naive model that predicts the last observed value can achieve competitive MSE but exhibits near-zero correlation with true market trends, making it ineffective for practical trading strategies. These shortcomings underscore the need for financial-specific benchmarks that reflect the complexity of real-world markets, along with evaluation frameworks that assess both predictive accuracy and economic utility.
@@ 2.1 Time-series Forecasting Models
Recent advances in time-series forecasting have focused on enhancing model capabilities through multi-scale decomposition, attention mechanism optimization, lightweight architectures, and representation learning innovations. Decomposition-based methods remain foundational for handling non-stationary signals. For example, Autoformer Wu et al. (2021a) pioneers autocorrelation-driven periodicity detection, replacing traditional moving averages with adaptive decomposition, while FEDformer Zhou et al. (2022) leverages Fourier-Wavelet hybrid spectral analysis to achieve multi-resolution frequency decomposition. Building on this, Non-stationary Transformers Liu et al. (2022) introduce dynamic stationarization modules that learn to normalize non-stationary inputs before applying attention mechanisms, significantly improving robustness to distribution shifts. These approaches highlight the synergy between frequency-domain analysis and temporal modeling.
@@
To address computational bottlenecks in Transformers, researchers have developed sparse attention variants: Informer Zhou et al. (2021) reduces quadratic complexity via probabilistic sparse attention with KL-divergence-based token selection, while Crossformer Zhang & Yan (2023) designs hierarchical cross-resolution attention to capture dependencies across temporal scales. Koopa Liu et al. (2023a) bypasses attention entirely, proposing time-varying Koopman dynamic systems that model latent state evolution through linear operators. Concurrently, lightweight architectures challenge conventional wisdom-DLinear Zeng et al. (2023) demonstrates that decoupled linear projections for trend and residual components can outperform complex models, while TSMixer Chen et al. (2023) and TimeMixer Wang et al. (2024a) adopt pure MLP-based architectures and hybrid frequency-time operators, respectively, to balance accuracy and computational cost.
@@
Representation learning breakthroughs further expand modeling capabilities. TimesNet Wu et al. (2023) transforms 1D sequences into 2D temporal matrices via period-phase folding, enabling 2D convolutions to capture intra-period and inter-period patterns simultaneously. PatchTST Nie et al. (2023) introduces channel-independent patching inspired by vision transformers, learning localized temporal embeddings through overlapping segments. iTransformer Liu et al. (2024) inverts the traditional architecture by treating variates as tokens and time points as features, enhancing multivariate dependency modeling. MICN Wang et al. (2023) employs multi-scale dilated convolutional pyramids to extract local periodic features hierarchically. Complementary techniques like RevIN Kim et al. (2021) address distribution shifts via bidirectional instance normalization, and TiDE Das et al. (2023) integrates temporal encoding with dense residual connections for efficient long-horizon forecasting. These innovations collectively advance three core principles: 1) Hybridization of frequency-temporal analysis, 2) Strategic simplification guided by signal processing theory, and 3) Systematic robustness enhancement through normalization and distribution alignment.
@@ 2.2 Financial task-related Studies
Recent advances in financial time series forecasting have increasingly adopted multi-horizon prediction frameworks to capture evolving market dynamics. Unlike traditional approaches that focus on single-step forecasting (e.g., predicting next-step prices or up-down trends as in Ding et al. (2015); Abe & Nakayama (2018); Kraus & Feuerriegel (2017); Sun et al. (2023); Li et al. (2024)), recent time series methods like Liu et al. (2024); Wu et al. (2023); Liu et al. (2023a) employ sequence-to-sequence architectures to predict multi-step trajectories of exchange time series, which helps to understand future temporal dynamics but also introduces challenges for prediction. Although large models and Reinforcement Learning methods are becoming increasingly popular in financial time series forecasting Nie et al. (2024); Li et al. (2024); Zong et al. (2024), the potential of small models and supervised learning approaches in financial time series forecasting has not been fully explored. It is important to build a bridge between cutting-edge models and financial time series tasks.
@@ 2.3 The Predictability of Asset Prices
Stock price prediction is indeed challenging, but it is not entirely impossible. This is evidenced by the long-standing work of quantitative hedge fund institutions, which continuously mine predictive indicators and patterns from historical datasets to cope with the ever-changing market environment, achieving excess returns. Moreover, some studies based on the Efficient Market Hypothesis (EMH) suggest that the stock market is not necessarily in the semi-strong form or strong-form efficiency as previously thought, which would prevent future price movements from being forecasted based on available historical information Miller et al. (1970); Malkiel (2003). On the contrary, many markets often fall between semi-strong and weak-form efficiency Lo & MacKinlay (1988), and Efficiency (1993) has validated the momentum effect in the stock market, where stocks that performed well in the past may continue to perform well in the future, thereby challenging the weak-form efficient market hypothesis. Fama & French (1996) discusses the explanatory power of multi-factor models for asset pricing anomalies, indirectly indicating the level of market efficiency. Additionally, advancements in technology have significantly enhanced data processing capabilities, leading to improvements in the performance of predictive models. This development brings data processing power and predictive performance that were previously unavailable in models based solely on historical data Brogaard & Zareei (2023); Leippold et al. (2022); Gu et al. (2020); Feng et al. (2020); Barberis et al. (2005).
@@ 3 Dataset Curation
Current research on long-term time series forecasting predominantly focuses on eight mainstream datasets Wu et al. (2023); Chen et al. (2023); Liu et al. (2024; 2023a); Zeng et al. (2023); Nie et al. (2023). Among these, five are related to electricity, while the remaining three pertain to weather, traffic, and exchange rates. Although exchange rate data falls under the financial domain, it is often overlooked compared to the other seven datasets or replaced by the ILI dataset from the illness area, due to the non-stationary and non-periodic characteristics of exchange rate data Liu et al. (2023b), which makes it more challenging to predict than other time series data that exhibit significant periodicity and stationarity.
While some state-of-the-art time series models have demonstrated strong long-term forecasting capabilities on mainstream datasets, they lack the robustness needed to tackle the complexities of real-world time series problems Tan et al. (2024); Bergmeir (2024a; b; c). To better investigate these real-world challenges, we propose three financial time series datasets.
@@ 3.1 Data Sources
We have constructed three financial time series datasets: GSMI, OPTION, and BTCF, each representing different subfields of finance. The GSMI dataset includes 20 major indices from global stock markets, recording daily price and trading volume data for these indices over nearly 20 years, from 2005 to 2024. The OPTION dataset includes the CSI 300ETF options from the Chinese financial market, with variables related to risk for both call and put options. The BTCF dataset contains hourly frequency data for Bitcoin spot and perpetual contracts, helping to understand the spot-contract lag relationship and facilitate long-short trading strategies Bulun (2020); Narayanasamy et al. (2023). In Table 3, we present the statistical properties of these three financial time series datasets with more details.
@@ 3.2 Data Preprocessing Methods
For the trading volume series $V_{0,t}=\{v_{0},v_{1},v_{2},...,v_{i},...,v_{t}\}$, the logarithmic transformation method is
@@ 4 New Evaluation Metrics
While mainstream time series forecasting works in top conferences often adopt error-based methods, such as MSE and MAE, as evaluation metrics for model predictions, these metrics face significant challenges when applied to financial time series. A simple example is that a Naive model as in Table 13, which directly uses the last observed value of the input sequence as the forecast, achieves remarkably low prediction errors on the Exchange dataset (we provide detailed analysis in Section B.5). This raises critical questions about the evaluation of financial time series forecasting, suggesting that correlation-based metrics should be introduced alongside prediction errors, as they are crucial for real-world financial applications. Although traditional Information Coefficient and Information Ratio metrics Treynor & Black (1973); Grinold & Kahn (2000) can measure temporal correlations in single-step univariate forecasting, they fail to assess multi-variable multi-step predictions. To address this limitation, we propose multi-step IC and multi-step IR, abbreviated as msIC and msIR, respectively.
@@ 5 Experiment
To bridge the gap between real-world financial time series data and cutting-edge time series models, we employ over 10 advanced time series models and conduct extensive experimental tests across three financial time series scenarios: Multivariate-to-Multivariate Forecasting, Multivariate-to-Univariate Forecasting, and Multivariate-to-Partial Forecasting. These time series models include: TimeMixerWang et al. (2024a), KoopaLiu et al. (2023a), iTransformerLiu et al. (2024), PSformerWang et al. (2024b), TiDEDas et al. (2023), PatchTSTNie et al. (2023), DLinearZeng et al. (2023), StationaryLiu et al. (2022), TSMixerChen et al. (2023), TimesNetWu et al. (2023), FEDformerZhou et al. (2022), AutoformerWu et al. (2021a), CrossformerZhang & Yan (2023), TransformerVaswani (2017), InformerZhou et al. (2021), and a Naive model.
@@ 5.3 Multivariate-to-Partial Forecasting
Setup. Multivariate prediction of partial variables is not currently the mainstream experimental setup in time series forecasting. However, Wang et al. (2024d) discusses the importance of this practical scenario. In this work, we set the closing prices of 20 indices in the GSMI dataset as the target variables for prediction. We evaluate the performance of the models on the GSMI dataset. Additionally, we construct a portfolio selection strategy and backtest the return performance while holding different numbers of indices simultaneously.